# Partially Stochastic Infinitely Deep Bayesian Neural Networks

### v2

@@ 6 Conclusion and Future Work
We introduce Partially Stochastic Bayesian Neural Networks in the infinite-depth limit. We provide an exact mathematical formulation of our method and conceptualise how to split weights of continuous-time networks into subsets of stochastic and deterministic weights. With these foundations, we then prove several theoretical building blocks that lead to the conclusion that our novel family of networks are UCDAs. Additionally, we back up the results with a collection of empirical experiments that demonstrate how all our models outperform fully stochastic networks and other approximations of Bayesian inference in continuous-time models (e.g. MFVI). More importantly, we show that this improved performance is achieved with a significant efficiency boost that does not come with any performance trade-off. Furthermore, we provide a comprehensive study of the most important hyperparameters of our models. Therefore, our newly introduced family of Bayesian Neural Networks in the infinite-depth limit, are a better, more efficient, and practical alternative to existing models.
In future work, we aim to investigate models incorporating both horizontal and vertical cuts to explore potential performance improvements. Additionally, we will explore variance reduction techniques such as Sticking The Landing (Roeder et al., 2017) to improve the performance and stability of our models. Furthermore, an interesting direction to evaluate the generalization power of our approaches is to test them on a variety of datasets, including but not limited to Omniglot for few-shot learning and ImageNet for performance in high-resolution datasets. As scalability is a common issue with Bayesian models, implementing PSDE-BNN on learned features might help further improve the efficiency of the model. At last, it is relevant to investigate the relationship between the test set performance and the quality of the variational approximation more systematically, including exploring different priors and model complexities to better understand how the KL divergence impacts generalization across various settings.
@@ Acknowledgements
We thank the CDT of Mathematics in Random Systems: Analysis, Modelling, and Algorithms for backing the entirety of our research team. SCO and YS also extend gratitude to the Oxford-Man Institute of Quantitative Finance for their financial support and computing power. The manuscript has benefited from the insightful feedback of Harrison Waldon, Fayçal Drissi, and Jiajun He, for which all authors are appreciative. SCO would like to specifically acknowledge Richard Bergna for the stimulating discussions that sparked several of the ideas underpinning our research.
@@ Appendix C Further Training Visualisations
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption

### v3

@@ 6 Conclusion and Future Work
We have introduced Partially Stochastic Bayesian Neural Networks (PSDE-BNNs) in the infinite-depth limit, providing an exact mathematical formulation and demonstrating how to partition the weights of continuous-time networks into stochastic and deterministic subsets. Building on these foundations, we established several theoretical results that lead to the conclusion that our novel family of networks are Uniformly Continuous Depth Approximators (UCDAs). Our empirical experiments show that PSDE-BNNs outperform fully stochastic networks and other Bayesian inference approximations in continuous-time models, such as Mean Field Variational Inference (MFVI). Notably, this enhanced performance is achieved with a significant efficiency gain, without any trade-off in performance. Additionally, we conducted a thorough investigation of the key hyperparameters in our models. In summary, our newly introduced family of Bayesian Neural Networks in the infinite-depth limit offers a superior, more efficient, and practical alternative to existing models.
In future work, we aim to investigate models that incorporate both horizontal and vertical cuts to explore potential performance improvements. Additionally, we will explore variance reduction techniques such as Sticking The Landing (Roeder et al., 2017) to enhance the performance and stability of our models. Another promising direction is to evaluate the generalization capabilities of our approaches on a diverse range of datasets, including Omniglot for few-shot learning and ImageNet for high-resolution performance. Given that scalability is a common challenge with Bayesian models, implementing PSDE-BNN on learned features may further improve model efficiency. Finally, we plan to systematically investigate the relationship between test set performance and the quality of the variational approximation. This includes exploring different priors and model complexities to better understand how the KL divergence impacts generalization across various settings.
@@ Acknowledgements
We thank the CDT of Mathematics in Random Systems for backing the entirety of our research team. SCO and YS also extend gratitude to the Oxford-Man Institute of Quantitative Finance for their support. The manuscript has benefited from the insightful feedback of H. Waldon, F. Drissi, and J. He, for which all authors are appreciative. SCO would also like to thank R. Bergna for the stimulating discussions.